OpenStudy (anonymous):

Will medal. Jackson wants to put a fence around the perimeter of his rectangular yard, which is 10 1/2 feet 15 3/4 by feet. The fencing costs $5.40 per foot. How much will it cost Jackson to put a fence around the perimeter of his entire yard?

OpenStudy (anonymous):

so you know one side is 10.5 feet (1/2=0.5) and the other side is 15.75 (3/4=0.75), in order to find what the decimals would be cost wise use this to help: 1ft/5.40 =.5(or.75, whichever you want to do first but do both)/X find x for both and add them up with how much it is for the feet. hope this helped

OpenStudy (anonymous):

perimeter =52.5 times cost 5.4 =

OpenStudy (anonymous):

52.5 * 5.4 = 283.5 ?